OpenStudy (anonymous):

The equation, y = 10(1.0605)x, represents the winning prize money, in thousands, for first place in the Indianapolis 500 auto race. x represents the number of years after 1911, the first year. What is the value for 1911 (x = 0)? This is the prize money, in thousands. For instance, if you got 23, this represents $23,000. @mathstudent55

OpenStudy (anonymous):

@vinnv226

OpenStudy (anonymous):

explain step by step please, I have a lot of question like this

OpenStudy (mathstudent55):

If you use the given equation, you can find out how much money the winner of the race won as prize money. The equation is \(y = 10(1.0605)x \)

OpenStudy (mathstudent55):

The important thing here is to understand what \(x\), the variable means here. Before we get to \(x\), let's deal with \(y\) which is easier. After you input some number for \(x\) and calculate \(y\), that tells you what the prize money was.

OpenStudy (anonymous):

do you have a formula

OpenStudy (mathstudent55):

\(x\) represents how many years since 1911. For example, if you want to know how much the prize money was in 1912, since 1912 was 1 year from 1911, you'd evaluate the expression with \(x = 1\) since 1912 - 1911 = 1. If you want to know how much the prize money was in 1923, you'd do 1923 - 1911 = 10. That means you need to use \(x = 10\). In other words, to find the prize money in year k, first subtract k - 1911, then enter the result of that subtraction as x.

OpenStudy (anonymous):

okay I understand that!

OpenStudy (anonymous):

10*1.0605??

OpenStudy (mathstudent55):

Ok. The first question using the formula is what do you get as prize money for 1911? Well, 1911 is 1911 - 1911 = 0, 0 years after 1911, so evaluate the expression with x = 0. y = 10(1.0605)x y = 10(1.0605)(0) y = 0 What does this y = 0 value mean? You may be tempted to say that this means the prize money for 1911 is zero, but I don't think you can say that. This equation is defined for years after 1911. For example, you can use the formula for 1912, 1913, 1914, etc. This formula is not defined for 1911, so whatever value you get for 1911 is meaningless.